solve this: A is 3 times as old as B and B is 5 years older than C. If B is X years old write the algebraic expressions for following.
a) A's age today
b) C's age today
c) A's age in 6 years time
d) The sum of A and B age in 4 years time

Answer by Edwin McCravy(20056)

solve this: A is 3 times as old as B and B is 5 years older than C. If B is X years old write the algebraic expressions for following.
1.  A is 3 times as old as B 
    A = 3B


2.  B is 5 years older than C. 
    B = C+5

3.  B is X 
    B = X

a) A's age today
From (1),  A = 3B
From (3),  B = X
Substituting X for B
           A = 3X

---------------------------

b) C's age today
From (2), B = C+5
From (3), B = X
Substitute X for B
          X = C+5
Subtract 5 from both sides
        X-5 = C

---------------------------

c) A's age in 6 years time
Since from a), A = 3X  
Add 6 to both sides:
             A+6 = 3X+6    

Answer: 3X+6

---------------------------
d) The sum of A and B age in 4 years time
A is 3X now, so in four years, A will be 3X+4
B is  X now  so in four years, B will be X+4

In four years their sum will be

                (3X+4)+(X+4) = 3X+4+X+4 = 4X+8

Answer: 4X+8

Edwin